### v1

# Tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2

Abstract
Ni0.33NbSe2 has been identified as an altermagnet with a Néel temperature of 84 K, although Ni0.21NbSe2 behaves like a nonmagnetic metal. In this work, we explore the magnetic phases of Ni-intercalated NbSe2 with stoichiometry Ni0.25NbSe2 (also represented as NiNb4Se8), which lies between the two previously studied compositions. Our results reveal a transition in the ground state from a stripe antiferromagnetic phase with Kramers degeneracy (below the critical value of Coulomb repulsion UC) to a ferromagnetic phase (above UC). In the Kramers antiferromagnetic state, the nearest-neighbour magnetic coupling is ferromagnetic between the layers and forms stripe-like patterns in the ab plane. This behavior contrasts with that of Ni0.33NbSe2 where the nearest-neighbour magnetic couplings are opposite; this can be attributed to the different Wyckoff positions of the magnetic atoms, which produce different space groups and different magnetic space groups. Spin-orbit coupling (SOC) reduces the value of UC, and the magnetic easy axis aligns along the z-axis. The strong magnetocrystalline anisotropy and the second-nearest neighbor coupling stabilize the Kramers antiferromagnetic and ferromagnetic phases with spins aligned along the z-axis over the non-collinear 120∘ phase with in-plane spin orientations. Given the triangular lattice and the closeness to both ferromagnetic and antiferromagnetic coupling, a metamagnetic transition is possible. Furthermore, a spin-glass phase may emerge if the distribution of intercalated Ni atoms is spatially inhomogeneous. Substantial modifications to the electronic structure accompany Ni intercalation in NbSe2 with respect to the pristine phase. Compared to pristine NbSe2, which features a hole pocket at the point, the stripe phase of Ni0.25NbSe2 instead exhibits an electron pocket at . This pocket at has predominantly 3z2-r2 character, and derives its main spectral weight from the Nb atoms. Ni intercalation also shifts the Van Hove singularity away from the Fermi level, thereby suppressing potential electronic instabilities.
I Introduction
The tunability of magnetic and electronic phases in transition metal dichalcogenides (TMDs) doped or intercalated with ionic liquidsZhang et al. (2022), through self-intercalationWang et al. (2024), or with moleculesScheidt et al. (2015) has attracted significant research interest due to its potential for spintronic applications and the generation of new exotic phases. Intercalation of transition metals into TMDs modifies the electronic structure, leading to changes in the exchange interactions that govern the magnetic ordering. Focusing on the intercalation with magnetic transition metals, these materials exhibit diverse magnetic behaviors, including Weyl ferromagnetismSaika et al. (2022), chiral helimagnetismCao et al. (2020), altermagnetismVita et al. (2025) and spin glass statesKong et al. (2023); Maniv et al. (2021). This depends on the choice of dopant, concentration, and external factors such as strain and temperatureHe et al. (2024). The spatial distribution and local environment of the intercalate atoms plays a crucial role in defining the resultant magnetic order. The degree of hybridization between transition metal d-states and chalcogen p-states can significantly impact the strength and the nature of magnetic interactions.Liu et al. (2024) External stimuli, including electric gating, pressure, and strain engineering, offer further tunability. Strain can enhance SOC and modify superexchange interactions, leading to novel magnetic phases. Similarly, electrostatic gating has been demonstrated to control magnetic order dynamically, enabling voltage-driven phase transitions. These tunable properties make magnetic TMDs promising candidates for next-generation spintronic devices, including nonvolatile memories and logic applications.Nair et al. (2020); Shao et al. (2022)
Several transition metal dichalcogenides host electronic instabilities such as charge density wave (CDW)Skolimowski et al. (2019); Wang et al. (2020); Cossu et al. (2020) and superconductivityXiao et al. (2022); Lei et al. (2020). In the ultrathin limit, NbSe2 hosts superconductivityUgeda et al. (2016); Xi et al. (2016) and CDWRivano et al. (2025), which can interplay with different layered stacking ordersCossu et al. (2024); Wang et al. (2022), self-intercalationYin et al. (2023) and can host topological phasesChiu et al. (2025). One of the relevant ingredients for the electronic instabilities in NbSe2 is the presence of Van Hove singularity (VHS) close to the Fermi level.Calandra et al. (2009); Borisenko et al. (2011) Several papers reported the successful synthesis of NbSe2 intercalated with early 3d transition metalsHe et al. (2024), while very few results have been reported for Ni intercalation, which is the focus of the present paper. When intercalating Ni in NbSe2, the nearest neighbor (NN) magnetic exchange tends to be antiferromagnetic between the planes in the case of NiNb3S6An et al. (2023). NiNb3S6 is altermagnetic with a critical temperature of 84 KAn et al. (2023) and shares the same structure as the altermagnet VNb3S6Ray et al. (2025). The NiNb4Se8 compound contains less Ni than NiNb3S6 (there are 0.25 Ni per Nb atom versus 0.33 Ni per Nb atom), but we still expect the system to be magnetic. We expect a modest critical temperature, assuming a linear dependence on Ni composition and that the magnetism is relatively insensitive to whether S or Se is present. The Ni atoms are positioned on top of the Nb atoms. The distribution of the Ni atoms is assumed to be uniform in the Density functional theory (DFT) calculations, while in the experiment, there are vacancies that will add disorder to the magnetic structure. Recently, a compound with the same space group CoNb4Se8 was found to be altermagnetic due to the antiferromagnetic coupling between the Ni plane and antiferromagnetic coupling in the ab planeRegmi et al. (2025); Vita et al. (2025). The only theoretical study about NiNb4Se8 focuses on the small unit cell (two magnetic atoms) within standard DFT, finding that the system is ferromagnetic with a low magnetic momentRamakrishna et al. (2024). We will see that we reproduce these results under the same conditions, but we expand the study to all possible magnetic phases. Experimentally, a small amount of Ni intercalation in Ni0.21NbSe2 produces a nonmagnetic system and an electron pocket at the pointWadge et al. (2025), differently from the hole pocket of the pristine NbSe2. The strong modifications of the Fermi surface, by the Ni atoms with respect to the pristine NbSe2, disrupt the CDWWadge et al. (2025).
In this paper, we study the magnetic properties of NiNb4Se8, including collinear and non-collinear phases, discussing the tunability of the magnetic properties in Ni intercalated transition metal dichalcogenide NbSe2. We will demonstrate that the NiNb4Se8 transitions from a stripe antiferromagnetic phase to a ferromagnetic phase as a function of U, and discuss the possibility of a spin-glass phase. We will highlight the difference with the altermagnetic NiNb3Se6 and estimate the magnetic exchanges. The paper is structured as follows: Section 2 discusses magnetic properties with SOC, while Section 3 investigates magnetic properties including relativistic effects. In the fourth and final Section, we draw our conclusions and outlook.
Atom | WP | |||
Ni | 0 | 0 | 0 | |
Nb(1) | 0 | 0 | ||
Nb(2) | 0.01949 | 0.50975 | ||
Se(1) | 0.83145 | 0.66290 | 0.88328 | |
Se(2) | 0.39321 | |||
II Electronic and Magnetic properties without SOC
In this section, we report the structural properties and the study of the collinear magnetic phases as a function of the electronic correlations for NiNb4Se8.
II.1 Structural relaxation and magnetic ground state
As it was confirmed from experimental studies for CoNb4Se8.Regmi et al. (2025), the position of the magnetic atoms tends to be on the top of Nb in transition metals intercalated NbSe2 with NiNb4Se8 stoichiometry. The experimental lattice constants of pristine bulk NbSe2 were 3.44445 Å and 12.54527 Å. The experimental lattice constants for the equivalent cell in NiNb4Se8 are 3.4540 Å and 12.428 Å, with a reduction of the in-plane lattice constant and an increase of the out-of-plane lattice constant.Wadge et al. (2025) The unit cell of NiNb4Se8 can be obtained from NbSe2 by doubling along the a and b-axis using a 221 supercell; this supercell contains two intercalated Ni atoms and allows us to simulate the magnetic phases in Fig. 1(a) and 1(b). By using a 421 supercell containing 4 magnetic atoms, we can simulate the magnetic phases reported in Fig. 1(c) and 1(d), which have two Ni atoms per plane. The Ni atoms are on a triangular lattice in the ab plane, while they are on top of each other regarding the out-of-plane connectivity. In all these 3d-intercalated transition metal dichalcogenides, we have both transition metals surrounded by 6 chalcogen atoms forming octahedra.
We use the unit cell to simulate the FM and AM phases, while we use the supercell for the AFM1 and AFM2 phases. We report the Brillouin zone (BZ) for the two space groups considered for this compound: P63/ (no. 194) for the unit cell and P21 (no. 4) for the supercell. We perform structural relaxation, constraining the volume to the experimental values. The outputs for the internal degrees of freedom are reported for the nonmagnetic phase (NM), the FM and the AM phase in Table 1, 2 and 3, respectively. We note that the internal degrees of freedom barely change from one magnetic phase to another.
Atom | WP | |||
Ni | 0 | 0 | 0 | |
Nb(1) | 0 | 0 | ||
Nb(2) | 0.01935 | 0.50968 | ||
Se(1) | 0.83135 | 0.66271 | 0.88330 | |
Se(2) | 0.39298 | |||
Atom | WP | |||
Ni | 0 | 0 | 0 | |
Nb(1) | 0 | 0 | ||
Nb(2) | 0.01967 | 0.50984 | ||
Se(1) | 0.83092 | 0.66185 | 0.88331 | |
Se(2) | 0.39262 | |||
Given that the supercell contains 4 atoms, we can consider the ferromagnetic phase (FM) and 3 independent magnetic phases with zero magnetization. The real space pictures of these magnetic phases are reported in Fig. 1. The FM phase is reported in Fig. 1(a). We will demonstrate that the magnetic phase with in-plane ferromagnetic coupling and out-of-plane antiferromagnetic coupling is altermagnetic, we name it AM, and we report it in Fig. 1(b). The other two magnetic configurations are antiferromagnetic with Kramers degeneracy; they are illustrated in Fig. 1(c-d) and we define them as AFM1 and AFM2, respectively. AFM1 hosts both in-plane and out-of-plane antiferromagnetic coupling, while AFM2 hosts ferromagnetic coupling out-of-plane and antiferromagnetic coupling. Within the single layer, AFM1 and AFM2 have the same magnetic order, which is usually named as stripe phase magnetic order. We present in Fig. 2 the energy differences between these 4 magnetic configurations. In our theoretical results, even for U=0, the ground state is magnetic. Therefore, we expect the system to be experimentally magnetic at low temperatures in the case of nominal stoichiometry and ideal crystal structure. Note that the AM phase results in Fig. 2 are reported starting from U=1 eV, since at U=0 eV, the AM phase tends to converge to a nonmagnetic state. As shown in Fig. 2, the FM and AFM2 configurations are the most stable phases. The ground state depends on the value of the electronic correlations, there is a critical value UC at which we have a change of the ground state. The AFM2 phase is the ground state for UUC while the ferromagnetic phase is the ground state for UUC. In the non-relativistic case, we estimate U2.1 eV. As we will see in the next sections, the system is metallic with a large density of states at the Fermi level. The metallicity strongly screens the Coulomb repulsion, lowering its value. Since the experimental magnetic moments of metallic Ni1/3NbS2 are reproduced using GGA+U with U1 eVAn et al. (2023); Tenzin et al. (2025), we infer that AFM2 is the likely ground state of NiNb4Se8 under ideal stoichiometry and ordered Ni distribution. The out-of-plane ferromagnetic coupling between the Ni atoms in Ni0.25NbSe2 differs from the out-of-plane antiferromagnetic coupling found in Ni0.33NbS2 despite the two-dimensional material hosts being isoelectronic. The difference lies in the connectivity between the Ni atoms. The Ni atoms of different layers in Ni0.25NbSe2 are one on top of each other as reported in Fig. 3(b), while they are not in Ni0.33NbS2 as we can see in Fig. 3(a).
II.2 Band structure
We calculate the band structure for the 4 magnetic phases and the nonmagnetic phase. The k-paths for the two space groups are reported in Fig. 4, while the band structures are reported in Fig. 5. We plot the results of the nonmagnetic phase at U=0 in Fig. 5(a), while we plot in Fig. 5(c) the band structure of the ferromagnetic phase for U=3 eV when it is the ground state. For the other phase, we plot the band structures at U=2 eV in Fig. 5(b,d,e).
We observe the non-relativistic spin-splitting in the FM and AM phases, while AFM1 and AFM2 host Kramers degeneracy. The AM band structure in Fig. 5(b) exhibits a non-relativistic spin-splitting up to 0.2 eV along L1--L2 with a g-wave spin-momentum locking. While the relaxed atomic positions are basically not affected by the magnetism, the band structure strongly changes across the magnetic phases. For instance, we can focus on the point. In bulk pristine NbSe2, at the point, there are bands with negative effective masses (hole pockets) in the low-energy band structureRossnagel et al. (2001)Rahn et al. (2012)Johannes et al. (2006). Unlike pristine NbSe2, all the magnetic phases feature an electron pocket with positive effective mass at the point in its low-energy band structure. The Ni intercalation changed the character of the carriers at the point from hole to electrons. To gain further insight, we focus on AFM2, which is the ground state at low Coulomb repulsion. We examined the orbital character of the electron band in the low-energy range at the point for the AFM2 phase. We find that the electron band below the Fermi level is mainly composed of the 3z2-r2 electrons of Nb with additional Nb components of xy and x2-y2 plus contribution from the same orbitals on the Ni atoms. In the FM phase, the electron pocket is composed of the same orbitals but only from Nb atoms.
II.3 Fermi surface
We report the Fermi surfaces in the first BZ for the AFM2 and FM phases in Fig. 6. To compare the two phases, we use for both cases a 421 supercell (see Appendix). In this particular supercell, we do not find the hexagonal BZ of the pristine NbSe2, but the BZ is a parallelepiped. The AFM2 phase hosts 5 Fermi surfaces as shown in Fig. 6(a). The Fermi surfaces for spin-up and spin-down channels of the FM phase are plotted in Fig. 6(a) and 6(a), respectively. There are 4 Fermi surfaces for the spin-up and 6 for the spin-down channel. In all phases, we have hole- and electron-pockets.
Unlike NbSe2, which exhibits doubly nested hole pockets around the pointRossnagel et al. (2001)Johannes et al. (2006), significant electron pockets emerge in both the AFM2 and FM phases upon Ni intercalation, in agreement with band structure calculations. Along with hole-pockets present in all figures, the electron pockets at are visible mainly in the third panel of Fig. 6(a), in the second panel of Fig. 6(b), and in the fifth panel Fig. 6(c). The characteristic isolated triangular pockets observed in pristine NbSe2 vanish, and in certain bands, the Fermi surface evolves into a fully complex interconnected topology. Furthermore, compared to NbSe2, the Fermi surface dispersion along the direction is markedly enhanced. The latter is due to the Ni atoms, which work as bridges between the NbSe2 layers, making the system more three-dimensional.
II.4 Nearest-neighbor spin exchange coupling
The nearest-neighbor spin exchange interaction parameter can be determined by mapping the system onto the Heisenberg model with effective spin S=1. The corresponding spin Hamiltonian can be expressed asWhite (2006):
Here, represents the total spin vector of the Ni atoms. For NiNb4Se8, Ni exists in an electronic configuration close to the Ni3+(3d7) state. This means that the triply degenerate t2g states are fully occupied, while the doubly degenerate eg states accommodate a single electron. Consequently, Ni adopts a low-spin state with a total spin of =. The magnetic moments extracted from the ground state phase AFM2 range from 0.436 at U=0 to 0.975 at U=3 eV, in agreement with an = electronic state.
We aim to calculate the nearest-neighbour magnetic exchange interactions: one in the out-of-plane direction and one in-plane. Considering the nearest-neighbor exchange interactions, we define the out-of-plane exchange constant as and the in-plane exchange constant as . To simplify the calculations, we map the Heisenberg model on spins with =1. In this framework, the total energies of these magnetic structures can be expressed as follows:
While is the nonmagnetic total energy of the system. The same equations are valid without and with SOC. To determine the nearest-neighbor exchange parameter from DFT calculations, we employ the Heisenberg Hamiltonian and establish a relationship between the total energies of just three distinct magnetic configurations: From the previous set of equations, we will consider the first three equations (2-4) to derive J∥ and J⟂. These equations will be used in the next Section to derive the magnetic exchanges J∥ and J⟂ in the non-relativistic and relativistic cases.
III Electronic and magnetic properties with SOC: role of the MCA
In this section, we report the study of the magnetic phases as a function of the electronic correlations for NiNb4Se8 in the relativistic case. With respect to the previous section, we include the investigation of the magnetocrystalline anisotropy and the study of the non-collinear 120∘ phase. Indeed, given the triangular lattice in the plane, if the system is antiferromagnetic in the plane, the magnetic system could become frustrated and the magnetism will become noncollinearDey and Botana (2020).
III.1 Magnetocrystalline anisotropy of the ferromagnetic phase
We calculate the magnetocrystalline anisotropy energy for the ferromagnetic phase at U=0 and U=3 eV. The easy axis lies along the z-axis. We can write the magnetocrystalline anisotropy as
where is the polar angle of the spinsBlundell (2001). We obtain that the magnetocrystalline anisotropy in the ferromagnetic phase is K2=0.6 meV per atom for U=0 and K2=0.9 meV per atom for U=3 eV. This relatively large value arises from the system’s strong crystalline anisotropy due to its layered crystal structureBaranov et al. (2011).
III.2 Magnetic ground state with SOC: NN magnetic couplings
Using the same strategy used in the previous section, we have calculated the energy differences of the four collinear magnetic phases in the presence of SOC. The results are reported in Fig. 7. Comparing these results with the ones without SOC, we obtain a reduction of UC to a value around 1.1 eV. For the rest, the calculation results reveal that SOC induces minor changes in the ground state energies of the four magnetic systems, though the overall trend remains largely unchanged. FM and AFM2 remain the most stable configurations. As in the case without SOC, the energy difference between the AFM1 and NM states is nearly identical at U=0, likely due to the small magnetic moment (0.27 with SOC). This will lead to a small compared to the results for other values of U.
Employing self-consistent energy differences, we calculate the magnetic exchanges and as defined in the previous sections for both relativistic and non-relativistic calculations. The structure used in the calculation was obtained from optimization under the AFM2 magnetic configuration. We compare the value of exchange constants and as a function of U, with and without SOC. Our exchange energy calculations provide a comprehensive explanation for the magnetic ground state energetics. Disregarding the small variation in U and the effects of SOC, the absolute value of is one order of magnitude greater than that of , demonstrating that the out-of-plane exchange energy is nearly significantly higher than the in-plane exchange energy. Therefore, we have a strongly anisotropic magnetism. From the perspective of SOC, it has a slight effect on the in-plane exchange energy , making the ferromagnetic state more favored as the ground state. The significantly negative out-of-plane exchange interactions not only strongly favor FM and AFM2 configurations with identical out-of-plane spin orientations as the dominant ground states, but also reduce the structural stability of the out-of-plane antiparallel spin magnetic configuration. Leading to an increase in the ground state energy of both the AM and AFM1 phases and even a tendency towards the nonmagnetic state. Furthermore, the in-plane exchange interactions are negligibly small, resulting in minimal energy differences between FM and AFM2 configurations with opposite in-plane spin alignments. The influence of U is minimal on , but it significantly affects , even forcing it to change sign, which further alters the magnetic ground state. For low values of U, the positive suggests that the AFM2 phase should be the ground state. However, as U increases, decreases notably, shifting the ground state from AFM2 to FM. Therefore, as previously mentioned, it is challenging to determine whether the ground state is FM or AFM2 based on the existing calculations, as U values between 1 and 2 eV are reasonable for NiNb4Se8.
DFT+U | DFT+U+SOC | |||
U () | ||||
0 | -2.79 | 0.76 | -1.99 | 0.31 |
1 | -7.88 | 0.90 | -6.58 | 0.09 |
2 | -9.26 | 0.03 | -7.73 | -1.10 |
3 | -9.11 | -1.08 | -7.16 | -2.44 |
III.3 Fermiology of the AFM2 and FM phase
We plot the relativistic band structure focusing on the most relevant magnetic phases, which are the FM phase and the AFM2 phase. Their band structures for U=2 eV are plotted in Fig. 8(a) and 8(b), respectively. As in the case without SOC, significant metallicity is found in both cases. In both cases, there are electron bands at the point close to the Fermi level, which is the opposite of the pristine case. The AFM2 phase hosts the Kramers degeneracy.
We plot the total density of states (TDOS) for the FM and AFM2 phases in Fig. 9(a) and 9(b), respectively, since these phases are the ground states in different regions of the phase diagram. In both cases, there is a significant DOS at the Fermi level, which produces the screening of the Coulomb repulsion, lowering the electronic correlations. We compare these DOS results with the pristine NbSe2, in which there is a VHS at -0.2 eV with respect to the Fermi level, particularly due to saddle points in the electronic band structure near the M point in the Brillouin zoneCalandra et al. (2009). The intercalated transition metal d-orbitals bands are narrower than the Nb bands and closer to the Fermi level. Once the Nb and Ni bands hybridize, the Nb bands are pushed away from the Fermi level. The intercalation of the Ni atoms shifts the VHS at -0.4 eV from the Fermi level, as we can see in both Figs. 9(a) and 9(b). As a result, this shift of the VHS reduces the TDOS at the Fermi level and can suppress the electronic instabilities, such as the CDW and superconductivityYan et al. (2019). It was shown that there is a strong interplay between a VHS and a CDW state in the isostructural and isoelectronic 2H-TaSe2Luckin et al. (2024), therefore, we can expect that Ni intercalation would affect the CDW. Indeed, after intercalation, the change of the Fermi surface stabilizes the lattice. On the other hand, Ni electrons are involved in the Nb-Se covalent bonds within each monolayer, which weakens the Nb-Nb interactions and prevents the clustering effect of Nb atoms responsible for the CDW distortion. Thus, because of these two factors, the CDW phase at low temperature is expected to be suppressed in the intercalated NbSe2.Yin et al. (2023).
III.4 Non collinear phases
The magnetic coupling is antiferromagnetic within the plane, therefore, our system is an effective 2D triangular lattice with antiferromagnetic first-neighbor interaction for low U and ferromagnetic first-neighbor interaction for high U. The collinear antiferromagnetic phase investigated in the previous section is the stripe phase. From the literature, we know that the ground state of a 2D triangular system is the stripe when the second neighbor magnetic exchange is antiferromagnetic above a threshold which is around =0.14; while the ground state is non-collinear 120∘ phase when the second neighbor interaction is ferromagnetic or weakly antiferromagneticHu et al. (2015). Between these two phases, the theory predicts the quantum spin liquid phase for insulating systems. Since the system is metallic, it cannot host the quantum spin liquid phase, but the competition between the stripe and the non-collinear phases holds. We constructed a supercell to simulate the non-collinear 120∘ phase, which is reported in Fig. 10. We performed the structural relaxation on the non-collinear phase in the presence of SOC, and we obtained that the non-collinear phase is 15 meV per formula unit higher in energy than the stripe. Therefore, we can assume that the value of is antiferromagnetic and it is the key parameter to stabilize the stripe AFM2 phase. Another factor that assisted the magnetic coupling in stabilizing the stripe phase is the large magnetocrystalline anisotropy. Indeed, the stripe phase can have spins along the easy axis (z-axis), while the non-collinear 120∘ phase has in-plane spins.
IV Conclusions
We performed first-principles calculations for the Ni-intercalated Nb2Se4 with stoichiometry NiNb4Se8. Our results indicate that the magnetic properties of this system are highly tunable, capable of hosting a rich variety of magnetic phases.
NiNb4Se8 is at the border between a ferromagnetic state and a stripe-like antiferromagnetic phase with Kramers degeneracy (referred to as the AFM2 phase). This configuration is characterized by strong out-of-plane ferromagnetic coupling and relatively weak in-plane magnetic interactions. Among the magnetic structures considered, the AFM2 phase emerges as the ground state in the regime of low Coulomb repulsion, consistent with experimental conditions, where screening effects and a deviation from half-filling reduce the effective interaction strength. The system exhibits an easy axis along and a strong in-plane second-nearest-neighbor antiferromagnetic interaction. However, in the case of disordered Ni distribution, the fact that NiNb3S6 has an antiferromagnetic coupling on the triangular lattice in the plane out-of-plane will likely push the experimental samples into a regime of spin-glass. Given the geometrical frustration and proximity to both ferromagnetic and antiferromagnetic regimes, a metamagnetic transition under an applied magnetic field is also plausible. The critical temperature is expected to be below 84 K, which is the Néel temperature of NiNb3S6; a further reduction of the percentage of Ni will push the system into the nonmagnetic phaseWadge et al. (2025).
Ni intercalation significantly alters the electronic structure near the Fermi level and makes the electronic properties more three-dimensional with respect to pristine NbSe2. It shifts the VHS away from the Fermi energy, reducing the TDOS and suppressing electronic instabilities. Nevertheless, the system remains metallic, with a substantial DOS at the Fermi level that provides effective screening of Coulomb interactions. Notable changes are observed in the Fermi surface compared to pristine NbSe2, with multiple sheets appearing in both the FM and AFM2 phases. In the AFM2 and FM phases, there is a switch of the charge carriers at the point since an electron pocket appears at dominated by 3z2-r2 orbital character, with the primary atomic weight localized on the Nb atoms. We look forward to this magnetic system’s experimental realization, which offers a promising platform for exploring tunable magnetism and correlated electronic behavior.
Acknowledgements.
The authors acknowledge A. Wadge, L. Plucinski, A. Wiśniewski, L. K. Tenzin, J. Sławińska and R. Sattigeri for useful discussions. This research was supported by the ”MagTop” project (FENG.02.01-IP.05-0028/23) carried out within the ”International Research Agendas” programme of the Foundation for Polish Science, co-financed by the European Union under the European Funds for Smart Economy 2021-2027 (FENG). C.A. acknowledges support from PNRR MUR project PE0000023-NQSTI. We further acknowledge access to the computing facilities of the Interdisciplinary Center of Modeling at the University of Warsaw, Grant g91-1418, g91-1419, g96-1808 and g96-1809 for the availability of high-performance computing resources and support. We acknowledge the access to the computing facilities of the Poznan Supercomputing and Networking Center, Grants No. pl0267-01, pl0365-01 and pl0471-01.Appendix A Computational details
We conducted calculations using the Vienna Simulation Package (VASP)Kresse and Hafner (1993); Kresse and Furthmüller (1996) based on density functional theory (DFT) by the projector-augmented wave (PAW) methodKresse and Joubert (1999). The exchange-correlation functional was described using the generalized gradient approximation (GGA) proposed by Perdew, Burke, and Ernzerhof (PBE)Perdew et al. (1996). To investigate the relationship between the ground state and the on-site Coulomb interaction U in detail, the DFT U method was involvedLiechtenstein et al. (1995), U was varied from 0 to 3 eV. Hund’s coupling JH was set to 0.15U, a value deemed suitable for 3 transition metal compounds. The k-point grid was set to for the primitive cell and for the supercell, with a total energy convergence criterion of eV and a cutoff energy of 450 eV. To calculate the magnetocrystalline anisotropy, we use a large k-grid for the unit cell. NiNb4Se8 crystallizes in the space group P63/ (no. 194). We need to double the lattice parameters in the plane using a 2 2 1 supercell; the lattice parameters of the supercell determined experimentallyWadge et al. (2025) become ==2=6.8889 Å and =12.54527 Å, this supercell is used to study the FM and AM phase. We need to further double the cell and construct the supercell 4 2 1 to study the AFM1 and AFM2 magnetic configurations, this last supercell holds the space group of P21. Lattice parameters are kept fixed while atomic positions are optimized for each magnetic structure.